OpenStudy (anonymous):

You are given seven cards numbered 1 through 7. You shuffle the cards and deal them out on a table. What is the probability that the cards will be dealt in order: 1, 2, 3, 4, 5, 6, 7?

OpenStudy (kropot72):

The number of permutations of the seven cards is 7!. The probability of getting any one of these is 1/7!

OpenStudy (kropot72):

@sihelp Do you understand?

OpenStudy (accidentalaichan):

I agree with @kropot72. If you are given 7 cards, you have a 1/7 chance of having them pulled in the order given above.

OpenStudy (kropot72):

\[Required\ probability=\frac{1}{7\times6\times5\times4\times3\times2\times1}\]

OpenStudy (kropot72):

There are factorial 7 (7!) permutations of the 7 different cards. 7! = 7 * 6 * 5 * 4 * 3 * 2 * 1 = 5040 The probability of getting any particular one of these 5040 permutations is 1/5040.

OpenStudy (anonymous):

oh i just plugged it in it makes more since doing it that way

OpenStudy (kropot72):

@sihelp What did you plug it into?

OpenStudy (anonymous):

i used the calculator with scrap paper doing it the same way you just showed me and got the same answer

OpenStudy (kropot72):

k